f(x)=-2x^2+2x+1
the x coordinate of the vertex is?
The y coordinate of the vertex is?
the equation of the line of symmetry is?
the maximum/minimum of the f(x) is (simplified fraction)?

Answer by nerdybill(7384)

f(x)=-2x^2+2x+1
.
Axis of symmetry = -b/2a = -2/(2*(-2)) = -2/-4 = 1/2
Therefore, the equation of the line of symmetry is:
x = 1/2
.
Plug above back into:
f(x)=-2x^2+2x+1
f(1/2)=-2(1/2)^2+2(1/2)+1
f(1/2)=-2(1/4)+1+1
f(1/2)=-(1/2)+2
f(1/2)=-(1/2)+(4/2)
f(1/2)= 3/2
.
the x coordinate of the vertex is? 1/2
The y coordinate of the vertex is? 3/2
the equation of the line of symmetry is? x= 1/2
the MAXIMUM of the f(x) is (simplified fraction)? 3/2